x2y''-4xy'+4y=4x2-6x3

1
Expert's answer
2021-08-11T14:13:48-0400

Solution. Denote by D the linear differential operator "x^2\frac{d^2}{dx^2}-4x\frac{d}{dx}+4".

For all "k\in\mathbb{Z}" "D(x^k)=k(k-1)x^k-4kx^k+4x^k=(k^2-5k+4)x^k=(k-1)(k-4)x^k"

One can conclude, that the functions "x^k" are eigenfunctions for this operator and that the linear independent functions "x" and "x^4" belong to the kernel (null space) of "D".

Since "D" is an ordinary differential operator of the 2nd order, its kernel is 2-dimensional. Therefore, "\ker D={\rm span}\{x, x^4\}=\{Ax+Bx^4| A,B\in\mathbb{R}\}".

It remains to find any partial solution to form the general solution of the given ODE.

One can notice, that the right-hand function in this ODE lies in the span of eigenfunctions "x^2" and "x^3":

"D(c_1x^2+c_2x^3)=3c_1x^2-2c_2x^3=4x^2-6x^3". Therefore, "c_1=4\/3", "c_2=3", the function "\frac{4}{3}x^2+3x^3" is a partial solution of ODE, and hence, the general solution is

"y(x)=\frac{4}{3}x^2+3x^3+Ax+Bx^4"


Answer. "y(x)=\frac{4}{3}x^2+3x^3+Ax+Bx^4"